### v1

# Triple-top-gate technique for studying the strongly-interacting 2D electron systems in heterostructures

Abstract
We develop a technique that dramatically reduces the contact resistances and depletes a shunting channel between the contacts in the ungated area outside the Hall bar in ultra-high mobility SiGe/Si/SiGe heterostructures. It involves the creation of three overlapping independent gates deposited on top of the structure and allows transport measurements to be performed at millikelvin temperatures in the strongly interacting limit at low electron densities, where the energy of the electron-electron interactions dominates all other energy scales. This design allows one to observe the two-threshold voltage-current characteristics that are a signature for the collective depinning and sliding of the electron solid.
pacs:
71.10.Hf, 71.27.+a, 71.10.Ay
Ultra-high mobility Si/SiGe heterostructures have proved to be a handy tool for studying the effects of strong particle-particle correlations on the transport and thermodynamic properties of two-dimensional (2D) carrier systems. In particular, the fractional quantum Hall effect (see, e.g., Refs. Nelson et al. (1992); Lai et al. (2004a, b); Dolgopolov et al. (2021)), the 2D metal-insulator transition Coleridge et al. (1997); Lam et al. (1997); Melnikov et al. (2019), the band flattening Melnikov et al. (2017a); Dolgopolov et al. (2022), the metallic state in a spinless two-valley electron system Melnikov et al. (2020), strong enhancement of the effective mass and its independence of the electrons’ spins Melnikov et al. (2023), strong correlations and spin effects in transport Shashkin et al. (2020); Shashkin and Kravchenko (2021); Shashkin et al. (2022), and collective depinning and sliding of the quantum Wigner solid Melnikov et al. (2024) have been observed in this 2D carrier system. To study the strongly correlated regime, one must reach the limit where the Coulomb energy, $E_{\text{C}}$, greatly exceeds the Fermi energy, $E_{\text{F}}$. The strength of electron interactions is characterized by the interaction parameter $r_{s}$, defined as the ratio between the Coulomb and Fermi energies and equal to $r_{\text{s}}=g_{\text{v}}/(\pi n_{\text{s}})^{1/2}a_{\text{B}}$, where $g_{\text{v}}$ is the valley degeneracy, $n_{\text{s}}$ is the areal density of the electrons, and $a_{\text{B}}$ is the effective Bohr radius in a semiconductor. Therefore, the lower the electron density, the stronger the interactions. However, at low electron densities, the resistance of the contacts to the 2D layer of the electrons for Si-based systems dramatically increases at millikelvin temperatures, which makes the transport measurements extremely challenging. To overcome this obstacle, a so-called split-gate technique was developed Kruithof, Klapwijk, and Bakker (1991); Wang et al. (1992); Heemskerk and Klapwijk (1998). Gaps in the gate metallization were introduced, which split the gate into several parts to which the voltage can be applied to maintain a high electron density near the doped contact regions, whereas, in the main part of the sample, the electron density can be controlled independently. These gaps were narrow enough ($<100$ nm) for the given gate-oxide thickness to provide a smoothly descending electrostatic potential from the high-density part to the low-density part. On the other hand, independent control of the electron density in different parts of the sample can, in principle, be reached using a technique of overlapping gates (see, e.g., Refs. Angus et al. (2007); Lim et al. (2009); Willems van Beveren et al. (2010); Mittag et al. (2018, 2019); Iwakiri et al. (2022), where the overlapping gates separated by Al oxide as an insulator were used).
In this Letter, we report a novel triple-top-gate technique that dramatically reduces the contact resistances and depletes the shunting channel between the contacts in the ungated area of the samples outside the Hall bar in ultra-high mobility SiGe/Si/SiGe heterostructures. This allows transport measurements to be performed at millikelvin temperatures and extremely low electron densities. In contrast to the planar design of split gates used previously, we use a vertical design where different gates do not lie in the same plane and overlap, in which case, for the parallel-plate capacitor, the electron density in a part of the sample is controlled by the gate closest to this part. We measure voltage-current ($V$-$I$) characteristics in both double-gate and new triple-gate samples. In the double-gate samples, as the electron density is decreased, the single-threshold $V$-$I$ characteristics stop changing below a particular value of $n_{\text{s}}$, which indicates the presence of a shunting conduction channel. Therefore, the strongly interacting limit at low electron densities cannot be reached in such samples. In triple-gate samples, the shunting channel can be depleted, which allows access to the low electron densities and observation of the two-threshold $V$-$I$ characteristics that are a signature for the collective depinning and sliding of the electron solid.
Figure 1: (a) Schematics of the layer growth sequence and the cross-section of the triple-gate samples. (b) Schematic top view on the samples with three independent gates. The Hall-bar-shaped gate 1 is hidden and is shown by the dashed line in the central part of the sample. Contact gate 2 and additional gate 3 are above gate 1.
To create a sample, we used SiGe/Si/SiGe quantum wells grown in UHVCVD (for details, see Refs. Lu et al. (2009, 2010); Huang et al. (2012)). The low-temperature electron mobility in our samples reaches $\approx 200$ m2/Vs which is higher than that in SiGe-based structures used earlier Schäffler (1997); Lu et al. (2009, 2010); Huang et al. (2012). The $\approx 15$ nm wide silicon (001) quantum well was sandwiched between Si0.8Ge0.2 potential barriers (Fig. 1). Contacts to the 2D layer consisted of $\approx 300$ nm Au0.99Sb0.01 alloy deposited in a thermal evaporator and then annealed for 3–5 minutes at $440^{\circ}$ C in N2 atmosphere. The samples were patterned in Hall-bar shapes with the distance between the potential probes of 150 $\mu$m and width of 50 $\mu$m using photo-lithography. An $\approx 80$-nm-thick SiO layer was deposited on the surface of the wafer in a thermal evaporator, and an $\approx 40$-nm-thick Al gate was deposited on top of SiO. In addition, a $20$ nm thick layer of NiCr was deposited on top of Al for better adhesion of subsequent layers. After that, the contact gate was fabricated, for which the structure was covered by an $\approx 200$-nm-thick SiO layer, and an $\approx 60$-nm-thick aluminum gate was deposited on top of SiO. The contact gate allowed maintaining high electron density $\approx 2\times 10^{11}$ cm-2 near the contact regions regardless of the electron density in the main part of the sample. No additional doping was used, and the electron density was controlled by applying a positive dc voltage to the gate relative to the contacts. To improve the quality of contacts and increase the electron mobility, we used a saturating infra-red illumination on the samples Melnikov et al. (2015, 2017b). The resistances of the contacts were below $10$ kOhm. In addition to the described double-gate samples, we used triple-gate samples to deplete the shunting channel between the contacts in the ungated area outside the Hall bar, which can reveal itself at the lowest electron densities in the insulating regime. In this case, the samples were patterned in Hall-bar shapes with the distance between the potential probes of 100 $\mu$m and width of 50 $\mu$m using photo-lithography. The main NiCr/Al Hall-bar gate was made similarly with a distinction that an $\approx 200$-nm-thick SiO layer was evaporated and that the two-resist undercut technique was employed to obtain smooth edges of the evaporated films and increase the outcome of good samples. After that, the contact gate was fabricated, for which the structure was covered by an $\approx 150$-nm-thick SiO layer, and an $\approx 40$-nm-thick aluminum gate was deposited on top of SiO. An additional Al gate for depleting the shunting channel was fabricated simultaneously with the contact gate. Measurements were carried out in an Oxford TLM-400 dilution refrigerator. The voltage-current characteristics in the insulating regime were measured in the main part of the experiments for which a dc voltage was applied between the source and the nearest potential probe over a distance of $25$ $\mu$m. The current was measured by a current-voltage converter connected to a digital voltmeter. The voltage-current curves were a little asymmetric with respect to the reversal of the voltage; we used the negative part plotted versus the absolute value of voltage for convenience. The electron density was determined by Shubnikov-de Haas oscillations in the metallic regime using a standard four-terminal lock-in technique in a frequency range of 1–11 Hz at currents 0.04–4 nA in the linear regime of response. Two double-gate samples and four triple-gate samples were studied; the results obtained on all triple-gate samples were similar.
In Fig. 2, we show the $V$-$I$ characteristics measured in the insulating regime at low electron densities in double-gate samples. With increasing applied voltage, the current stays near zero up to some threshold voltage and then the current sharply increases, the threshold voltage increasing with decreasing electron density. However, these single-threshold $V$-$I$ characteristics stop changing below $n_{\text{s}}\approx 6\times 10^{9}$ cm-2. This indicates the presence of a shunting conduction channel in the ungated area outside the Hall bar that is obviously related to residual donor impurities in the SiGe/Si/SiGe heterostructures. Thus, the strongly interacting limit at low electron densities cannot be reached in such samples.
Figure 2: $V$-$I$ characteristics measured in double-gate samples at different electron densities (from left to right): 7.95, 6.98, 6.21, and 5.82×$\times 10^{9}$ cm-2. $T=30$ mK.
Figure 3: $V$-$I$ characteristics measured in triple-gate samples at different electron densities (from left to right): 6.27, 6.09, 5.91, and 5.82×$\times 10^{9}$ cm-2. $T=30$ mK. The inset shows $V$-$I$ characteristics for $n_{\text{s}}=5.91\times 10^{9}$ cm-2 and $5.82\times 10^{9}$ cm-2 on an expanded scale. Arrows indicate the threshold voltages, $V_{\text{th}1}$ and $V_{\text{th}2}$, the dynamic threshold $V_{\text{d}}$ obtained by the extrapolation (dotted line) of the linear part of the $V$-$I$ curves to zero current, and the static threshold $V_{\text{s}}=V_{\text{th}2}$. The dashed lines are fits to the data; see text.
Applying a negative voltage to the additional gate in triple-gate samples depletes the shunting channel and reveals quite a different behavior of the $V$-$I$ characteristics with decreasing electron density, as shown in Fig. 3. The observed $V$-$I$ characteristics change with decreasing $n_{\text{s}}$ in the entire range of electron densities studied. At the lowest electron densities, a second threshold voltage arises on the $V$-$I$ curves. Namely, with increasing applied voltage, the current stays near zero up to the first threshold voltage, $V_{\text{th}1}$, and then the current sharply increases until the second threshold voltage, $V_{\text{th}2}$, is reached. Above the second threshold voltage, the slope of the $V$-$I$ curves is reduced, and the behavior becomes linear but not ohmic. These two-threshold $V$-$I$ characteristics are similar to those observed earlier in silicon metal-oxide-semiconductor field-effect transistors in Ref. Brussarski et al. (2018), which signals that the strongly interacting limit at low electron densities can be reached in the 2D electron system in SiGe/Si/SiGe heterostructures.
Within the phenomenological theory of pinned elastic structures, the observed two-threshold $V$-$I$ characteristics are interpreted as evidence for the collective depinning of an electron solid. The model of the collective depinning of the vortex lattice in type-II superconductors Blatter et al. (1994) was adapted for the case of the electron solid in Refs. Brussarski et al. (2018); Melnikov et al. (2024). As the applied voltage is increased, the depinning of the electron solid is indicated by the appearance of a current. Between the dynamic ($V_{\text{d}}$) and static ($V_{\text{s}}$) thresholds, the collective pinning of the electron solid occurs, and the transport is thermally activated: $I=\sigma_{0}\,(V-V_{\rm d})\,\exp[-U_{\rm c}(1-V/V_{\rm s})/k_{\rm B}T]$, where $U_{\rm c}$ is the maximal activation energy of the pinning centers, $\sigma_{0}$ is a coefficient, and $V_{\rm d}$ corresponds to the pinning force. At voltages exceeding the static threshold, the electron solid slides with friction over a pinning barrier, as determined by the balance of the electric, pinning, and friction forces, resulting in linear $V$-$I$ characteristics. The corresponding fits, shown by the dashed lines in Fig. 3, describe well the experimental two-threshold $V$-$I$ characteristics.
In summary, we have developed a triple-top-gate technique that dramatically reduces the contact resistances and excludes the influence of the shunting channel in the ungated area outside the Hall bar in ultra-high mobility SiGe/Si/SiGe heterostructures. Our vertical design with three overlapping independent gates allows transport measurements to be performed at millikelvin temperatures in the strongly interacting limit at low electron densities, where the energy of the electron-electron interactions dominates all other energy scales. In particular, this design allows one to observe the two-threshold $V$-$I$ characteristics that are a signature for the collective depinning and sliding of the electron solid.
Acknowledgements.
This investigation was supported by RSF (project no. 24-22-00122).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.